OpenStudy (theedwardsfamily):

HELP PLZ! Fast!!!! During his first year of life, Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form.

OpenStudy (unklerhaukus):

what is (1 - 40% ) of 12 3/4 lb ?

OpenStudy (theedwardsfamily):

um...idk

OpenStudy (theedwardsfamily):

how do i figure tht out?

OpenStudy (theedwardsfamily):

@iGreen.

OpenStudy (unklerhaukus):

1 - 40% = 100% - 40% = 60% = 60/100 = 6/10

OpenStudy (theedwardsfamily):

@Holly00d1248 @confluxepic @pooja195 @sleepyjess @TheSmartOne

OpenStudy (confluxepic):

Is this a test. @TheEdwardsFamily

OpenStudy (theedwardsfamily):

um...no just a unit checkpoint

OpenStudy (theedwardsfamily):

HELP!!!!!!!!!!!

OpenStudy (theedwardsfamily):

@SyedMohammed98

OpenStudy (theedwardsfamily):

ugh! nvrmind

TheSmartOne (thesmartone):

So you don't need help anymore?

OpenStudy (unklerhaukus):

6 /10 of 12 3/4 lb = 6/10 x 12 3/4 lb = (6 x 12 3/4 ) /10 lb = [(6 x 12) + (6 x 3/4 )] /10 lb =